# PRISM: A Robust Framework for Skill-based Meta-Reinforcement Learning with Noisy Demonstrations

### v2

@@ A.3 Meta-Test Phase
6:        Filter high-return trajectories $\tau_{\text{high}}^{i}$ from $\tau_{l}^{i}$ and $\tau_{\text{exp}}^{i}$ s.t. $G>\min_{\tau^{\prime}\in\mathcal{B}_{\text{on}}^{i}}G(\tau^{\prime})$.
7:        Store $\tau_{h}^{i}$, $\tau_{\text{exp}}^{i}$, and $\tau_{\text{high}}^{i}$ into $\mathcal{B}_{h}^{i}$, $\mathcal{B}_{\text{exp}}^{i}$, and $\mathcal{B}_{\text{on}}^{i}$.
8:     end for
9:     Compute prioritization factors: $P_{\mathcal{B}_{\text{off}}}$ and $\beta$.
10:     for gradient step do
11:        Update $\pi_{h,\theta}$, $Q_{h,\theta}$, $q_{e,\theta}$ using Eq. (A.6) with $\theta\leftarrow\theta-\lambda^{\text{lr}}_{h}\cdot\nabla_{\theta}(\mathcal{L}^{\text{critic}}_{h}(\theta)+\mathcal{L}^{\text{actor}}_{h}(\theta))$ (Meta-RL).
12:        Update $\pi_{\text{exp},\psi}$, $Q_{\text{exp},\psi}$ using Eq. (A.3) with $\psi\leftarrow\psi-\lambda^{\text{lr}}_{\text{exp}}\cdot\nabla_{\psi}(\mathcal{L}^{\text{critic}}_{\text{exp}}(\psi)+\mathcal{L}^{\text{actor}}_{\text{exp}}(\psi))$ (Skill exploration).
13:        Update $\pi_{l,\phi}$, $q_{\phi}$, $p_{\phi}$ using Eq. (A.5) with $\phi\leftarrow\phi-\lambda^{\text{lr}}_{l}\cdot\nabla_{\phi}\mathcal{L}_{\text{skill}}(\phi)$ (Skill refinement).
14:        Update reward model $\hat{R}_{\zeta}$ using Eq. (A.4) with $\zeta\leftarrow\zeta-\lambda^{\text{lr}}_{\text{reward}}\cdot\nabla_{\zeta}\mathcal{L}_{\text{reward}}(\zeta)$.
15:     end for
16:     if $k\bmod K_{\text{iter}}=0$ then
17:        Update $\bar{\pi}_{l,\phi}\leftarrow\pi_{l,\phi}$, $\bar{q}_{\phi}\leftarrow q_{\phi}$, and $\bar{p}_{\phi}\leftarrow p_{\phi}$.
18:        Reinitialize $\pi_{h,\theta}$.
19:     end if
20:  end for
Algorithm 2 PRISM: Meta-test phase
0:  Target task $\mathcal{T}$, high-level policy $\pi_{h,\theta}$, value function $Q_{h,\theta}$, task encoder $q_{e,\theta}$, and low-level policy $\pi_{l,\phi}$.
1:  Collect context $c^{\mathcal{T}}$ using $\pi_{h,\theta}$ with $\pi_{l,\phi}$, $e\sim\mathcal{N}(\mathrm{0},\mathrm{I})$.
2:  Compute task representation $e^{\mathcal{T}}\sim q_{e,\theta}(\cdot|c^{\mathcal{T}})$.
3:  for iteration $k=1,2,\dots$ do
4:     Collect high-level trajectory $\tau^{\mathcal{T}}_{h}$ using $\pi_{h,\theta}$ with $\pi_{l,\phi}$, $e^{\mathcal{T}}$.
5:     Store $\tau^{\mathcal{T}}_{h}$ into $\mathcal{B}^{\mathcal{T}}_{h}$.
6:     for gradient step do
@@
7:        Update $\pi_{h,\theta}$, $Q_{h,\theta}$ using Eq. (A.6) with $\theta\leftarrow\theta-\lambda^{\text{lr}}_{h}\cdot\nabla_{\theta}(\mathcal{L}^{\text{critic}}_{h}(\theta)+\mathcal{L}^{\text{actor}}_{h}(\theta))$
8:     end for
9:  end for
Appendix B Detailed Experimental Setup
@@
In this section, we provide a detailed description of our experimental setup. The implementation is built on PyTorch with CUDA 11.7, running on an AMD EPYC 7313 CPU with an NVIDIA GeForce RTX 3090 GPU. PRISM is implemented based on the official open-source code of SiMPL, available at https://github.com/namsan96/SiMPL. For the environment implementations, we used SiMPL’s code for the Kitchen and Maze2D environments, SkiLD’s open-source code for the Office environment at https://github.com/clvrai/skild, and D4RL’s open-source code for AntMaze at https://github.com/Farama-Foundation/D4RL/tree/master.
@@
The hyperparameters for low-level policy training were referenced from SPiRL (Pertsch et al., 2021). Additional details about the baseline algorithms are provided in Section B.1, while Section B.2 elaborates on the environments used for evaluation. Section B.3 explains the construction of offline datasets for varying noise levels, and Section B.4 details the network architectures and hyperparameter configurations for policies, value functions, and other models.
B.1 Other Baselines
@@
SAC (Haarnoja et al., 2018) is a reinforcement learning algorithm that incorporates entropy to improve exploration. Instead of a standard value function, SAC uses a soft value function that combines entropy, with the entropy coefficient adjusted automatically to maintain the target entropy. To enhance value function estimation, SAC employs double $Q$ learning, using two independent value functions. SAC learns tasks from scratch without utilizing meta-train tasks or offline datasets. For the Kitchen and Office environments, the discount factor $\gamma$ is set to $0.95$, while $\gamma=0.99$ is used for Maze2D and AntMaze environments.
@@
SAC+RND combines SAC with random network distillation (RND) (Burda et al., 2018), an intrinsic motivation technique, to enhance exploration. Like SAC, it learns tasks from scratch without meta-train tasks or offline datasets. RL hyperparameters are shared with SAC, and RND-specific hyperparameters are set to match those in PRISM. Additionally, for fair comparison, the ratio of extrinsic to intrinsic rewards is aligned with PRISM.
@@
PEARL (Rakelly et al., 2019) is a context-based meta-RL algorithm that leverages a task encoder $q_{e}$ to derive task representations, which are then used to train a meta-policy. PEARL adapts its learned policy quickly to unseen tasks without utilizing skills or offline datasets. Unlike the original PEARL, which does not fine-tune during the meta-test phase, we modified it to include fine-tuning on target tasks for a fair comparison.
@@
PEARL+RND extends PEARL by incorporating RND to enhance exploration. Like SAC+RND, the ratio of extrinsic to intrinsic rewards is set to match PRISM for fair comparison.
@@
SPiRL (Pertsch et al., 2021) is a skill-based RL algorithm that first learns a fixed low-level policy from an offline dataset and then trains a high-level policy for specific tasks. SPiRL’s loss function is detailed in Section 3, and for fair comparison, loss scaling is aligned with PRISM.
@@
SiMPL (Nam et al., 2022) is a skill-based meta-RL algorithm that uses both offline datasets and meta-train tasks. While it shares PRISM’s approach of extracting reusable skills and performing meta-train and meta-test phases, SiMPL fixes the skill model without further updates during meta-training. SiMPL’s loss function is also detailed in Section 3, and SiMPL’s implementation uses the same hyperparameters as PRISM to ensure a fair comparison.
B.2 Environmental Details
@@
Each subtask involves manipulating one object to its target state, while a full task requires sequentially completing four subtasks. The agent earns a reward of 1 for each completed subtask, with a maximum score of 4 achievable within a 280-timestep horizon. For instance, an example task, illustrated in Fig. B.1, requires the agent to complete the sequence: microwave $\rightarrow$ kettle $\rightarrow$ bottom burner $\rightarrow$ slide cabinet. The observation space is a 60-dimensional continuous vector representing object positions and robot state information, while the action space is a 9-dimensional continuous vector. Based on the task setup from Nam et al. (2022), we expanded the meta-train task set by adding two additional tasks, resulting in a total of 25 meta-train tasks and 10 meta-test tasks. Detailed task configurations are provided in Table B.1.
@@
Table B.1: List of meta-train tasks and meta-test tasks in Kitchen environment
@@
Figure B.2: Office environment: An example of performing a task ((shed2, drawer)$\rightarrow$(eraser1, container)$\rightarrow$(pepsi2, tray))
@@
An example task, depicted in Fig. B.2, requires the agent to sequentially complete: (shed2 $\rightarrow$ drawer), (eraser1 $\rightarrow$ container), and (pepsi2 $\rightarrow$ tray). The observation space is a 76-dimensional continuous vector, including object positions and robot state information, while the action space is an 8-dimensional continuous vector. The meta-train and meta-test sets include 25 and 10 tasks, respectively, similar to the configuration in the Kitchen environment. A detailed task list is provided in Table B.2.
@@
Table B.2: List of meta-train tasks and meta-test tasks in Office environment
@@
Figure B.3: Maze2D environment: Visualization of simulation and meta-training/task tasks in Maze2D
@@
Fig. B.3 (a) provides a visualization of the Maze2D environment, while Fig. B.3 (b) illustrates the meta-train and meta-test tasks. In Fig. B.3 (b), green squares indicate the goal points for 40 meta-train tasks, and red squares represent the goal points for 10 meta-test tasks. All tasks share the same starting point at (10, 10), marked by a blue cross. The observation space is a 4-dimensional continuous vector containing the ball’s position and velocity, while the action space is a 2-dimensional continuous vector.
@@
Figure B.4: AntMaze environment: Visualization of simulation and meta-training/task tasks in AntMaze
@@
Fig. B.4 (a) shows a simulation image of the AntMaze environment, and Fig. B.4 (b) depicts the meta-train and meta-test tasks. In Fig. B.4 (b), green squares mark the goal points for 20 meta-train tasks, while red squares denote the goal points for 10 meta-test tasks. All tasks share a common starting point at (5, 5), indicated by a blue cross. The observation space is a 29-dimensional continuous vector that includes the Ant’s state and its $(x,y)$ coordinates, while the action space is an 8-dimensional continuous vector.
B.3 Construction of Offline Dataset
@@
B.4 Hyperparameter Setup
In this section, we outline the hyperparameter setup for the proposed PRISM framework. For high-level policy training, we adopt the hyperparameters from SiMPL for the Kitchen and Maze2D environments. For the Office and AntMaze environments, we conduct hyperparameter sweeps using the Kitchen and Maze2D configurations as baselines.
Most of PRISM’s hyperparameters are inherited from SiMPL, given its multiple loss functions. However, parameter sweeps are specifically conducted for the proposed prioritized skill refinement and maximum return relabeling components. We explore prioritization temperature values $T\in[0.1,0.5,1.0,2.0]$ and KLD coefficients $\lambda^{\text{kld}}_{\text{exp}}\in[0,0.001,0.002,0.005]$ for skill exploration, selecting the best-performing configurations as defaults. Additionally, the ratio of intrinsic to extrinsic rewards is fixed at levels that show optimal performance in single-task SAC experiments.
For implementing the skill models ($\pi_{l},q,p$), we follow SPiRL by utilizing LSTM (Graves & Graves, 2012) for the skill encoder and MLP structures for the low-level policy and skill decoder. For implementing the high-level models ($\pi_{h},Q_{h},q_{e}$), we follow SiMPL by utilizing Set Transformer (Lee et al., 2019) for the task encoder and MLP structures for the high-level policy and value function. Additionally, for implementing the PRISM framework, we utilize MLP structures for $\pi_{\text{exp}}$, $Q_{\text{exp}}$, and $\hat{R}$. The detailed hidden network sizes are presented in Table B.3 and Table B.4. Table B.3 presents the network architectures (the number of nodes in fully connected layers) and the hyperparameters shared across all environments, while Table B.4 details the environment-specific hyperparameter setups.
Table B.3: Network Architecture and Shared Hyperparameters
Shared Hyperparameters Group Name Environments
@@ B.4 Hyperparameter Setup
High-level Discount Factor $\gamma_{h}$ 0.99
@@
Network size $\pi_{h},Q_{h}$ [128]$\times$6 [256]$\times$4 [128]$\times$6
Low-level Buffer size $\mathcal{B}^{i}_{\text{on}}$ 10K
@@
Network size $\hat{R}$ [128]$\times$3
Network size $\pi_{l}$ [128]$\times$6
Network size $p$ [128]$\times$7
@@
Exploration RND state dropout ratio 0.7
@@
Learning rate $\lambda^{\text{lr}}_{\text{exp}}$ 0.0003
Network size $\pi_{\text{exp}},Q_{\text{exp}}$ [256]$\times$4
Network size $f,\hat{f}$ [128]$\times$4
Table B.4: Environmental Hyperparameters
Environmental Hyperparameters Group Name Environments
@@
High-level Buffer size $\mathcal{B}^{i}_{h}$ 3000 3000 20000 20000
@@
Low-level Skill refinement $K_{\text{iter}}$ 2000 2000 1000 2000
@@
Exploration Buffer size $\mathcal{B}^{i}_{\text{exp}}$ 100K 200K 100K 300K
Discount factor $\gamma_{\text{exp}}$ 0.95 0.95 0.99 0.99
@@
Entropy coefficient $\lambda^{\text{ent}}_{\text{exp}}$ 0.2 0.2 0.1 0.1
KLD coefficient $\lambda^{\text{kld}}_{\text{exp}}$ 0.005 (Expert) 0.001 0.001 0.001
@@
Appendix C Additional Comparison Results
C.1 Performance Comparison
Fig. C.1 presents the learning curves of average returns for the algorithms on test tasks, corresponding to the experiments summarized in Table 1. Rows represent evaluation environments, and columns denote noise levels. PRISM consistently demonstrated superior robustness, outperforming all baselines across various environments and noise levels.

### v3

@@ B.3 Meta-Test Phase
7      Filter high-return trajectories $\tau_{\text{high}}^{i}$ from $\tau_{l}^{i}$ and $\tau_{\text{imp}}^{i}$ s.t. $G>\min_{\tau^{\prime}\in\mathcal{B}_{\text{on}}^{i}}G(\tau^{\prime})$.
8      Store $\tau_{h}^{i}$, $\tau_{\text{imp}}^{i}$, and $\tau_{\text{high}}^{i}$ into $\mathcal{B}_{h}^{i}$, $\mathcal{B}_{\text{imp}}^{i}$, and $\mathcal{B}_{\text{on}}^{i}$.
9
10  Compute prioritization factors: $P_{\mathcal{B}_{\text{off}}}$ and $\beta$.
11 for gradient step do
12      (Decoupled Policy Learning)
@@
      Update $\pi_{\text{imp},\psi}$, $Q_{\text{imp},\psi}$ using Eq. (B.3) with $\psi\leftarrow\psi-\lambda^{\text{lr}}_{\text{imp}}\cdot\nabla_{\psi}(\mathcal{L}^{\text{critic}}_{\text{imp}}(\psi)+\mathcal{L}^{\text{actor}}_{\text{imp}}(\psi))$.
      Update $\pi_{h,\theta}$, $Q_{h,\theta}$, $q_{e,\theta}$ using Eq. (B.4) with $\theta\leftarrow\theta-\lambda^{\text{lr}}_{h}\cdot\nabla_{\theta}(\mathcal{L}^{\text{critic}}_{h}(\theta)+\mathcal{L}^{\text{actor}}_{h}(\theta))$.
13      (Self-Improvement Skill Learning)
14      Update reward model $\hat{R}_{\zeta}$ using Eq. (B.5) with $\zeta\leftarrow\zeta-\lambda^{\text{lr}}_{\text{reward}}\cdot\nabla_{\zeta}\mathcal{L}_{\text{reward}}(\zeta)$.
      Update $\pi_{l,\phi}$, $q_{\phi}$, $p_{\phi}$ using Eq. (B.6) with $\phi\leftarrow\phi-\lambda^{\text{lr}}_{l}\cdot\nabla_{\phi}\mathcal{L}_{\text{skill}}(\phi)$.
15
16 if $k\bmod K_{\text{iter}}=0$ then
17      Update $\bar{\pi}_{l,\phi}\leftarrow\pi_{l,\phi}$, $\bar{q}_{\phi}\leftarrow q_{\phi}$, and $\bar{p}_{\phi}\leftarrow p_{\phi}$.
18      Reinitialize $\pi_{h,\theta}$.
19
20
Algorithm 1 SISL: Meta-Train Phase
1
Require: Target task $\mathcal{T}$, high-level policy $\pi_{h,\theta}$, value function $Q_{h,\theta}$, task encoder $q_{e,\theta}$, and low-level policy $\pi_{l,\phi}$.
2 Collect context $c^{\mathcal{T}}$ using $\pi_{h,\theta}$ with $\pi_{l,\phi}$, $e\sim\mathcal{N}(\mathrm{0},\mathrm{I})$.
3 Compute task representation $e^{\mathcal{T}}\sim q_{e,\theta}(\cdot|c^{\mathcal{T}})$.
4 for iteration $k=1,2,\dots$ do
5   Collect high-level trajectory $\tau^{\mathcal{T}}_{h}$ using $\pi_{h,\theta}$ with $\pi_{l,\phi}$, $e^{\mathcal{T}}$.
6   Store $\tau^{\mathcal{T}}_{h}$ into $\mathcal{B}^{\mathcal{T}}_{h}$.
7 for gradient step do
8      Update $\pi_{h,\theta}$, $Q_{h,\theta}$ using Eq. (B.4) with $\theta\leftarrow\theta-\lambda^{\text{lr}}_{h}\cdot\nabla_{\theta}(\mathcal{L}^{\text{critic}}_{h}(\theta)+\mathcal{L}^{\text{actor}}_{h}(\theta))$
9
10
Algorithm 2 SISL: Meta-test phase
Appendix C Detailed Experimental Setup
@@
In this section, we provide a detailed description of our experimental setup. The implementation is built on PyTorch with CUDA 11.7, running on an AMD EPYC 7313 CPU with an NVIDIA GeForce RTX 3090 GPU. SISL is implemented based on the official open-source code of SiMPL, available at https://github.com/namsan96/SiMPL. For the environment implementations, we used SiMPL’s code for the Kitchen and Maze2D environments, SkiLD’s open-source code for the Office environment at https://github.com/clvrai/skild, and D4RL’s open-source code for AntMaze at https://github.com/Farama-Foundation/D4RL/tree/master.
@@
The hyperparameters for low-level policy training were referenced from SPiRL (Pertsch et al., 2021). Additional details about the baseline algorithms are provided in Section C.1, while Section C.2 elaborates on the environments used for evaluation. Section C.3 explains the construction of offline datasets for varying noise levels, and Section C.4 details the network architectures and hyperparameter configurations for policies, value functions, and other models.
C.1 Other Baselines
@@
SAC (Haarnoja et al., 2018) is a reinforcement learning algorithm that incorporates entropy to improve exploration. Instead of a standard value function, SAC uses a soft value function that combines entropy, with the entropy coefficient adjusted automatically to maintain the target entropy. To enhance value function estimation, SAC employs double $Q$ learning, using two independent value functions. SAC learns tasks from scratch without utilizing meta-train tasks or offline datasets. For the Kitchen and Office environments, the discount factor $\gamma$ is set to $0.95$, while $\gamma=0.99$ is used for Maze2D and AntMaze environments. We utilize the open-source code of SAC at https://github.com/denisyarats/pytorch_sac.
@@
SAC+RND combines SAC with random network distillation (RND) (Burda et al., 2018), an intrinsic motivation technique, to enhance exploration. Like SAC, it learns tasks from scratch without meta-train tasks or offline datasets. RL hyperparameters are shared with SAC, and RND-specific hyperparameters are set to match those in SISL. Additionally, for fair comparison, the ratio of extrinsic to intrinsic rewards is aligned with SISL. We utilize the open-source code of RND at https://github.com/openai/random-network-distillation.
@@
PEARL (Rakelly et al., 2019) is a context-based meta-RL algorithm that leverages a task encoder $q_{e}$ to derive task representations, which are then used to train a meta-policy. PEARL adapts its learned policy quickly to unseen tasks without utilizing skills or offline datasets. Unlike the original PEARL, which does not fine-tune during the meta-test phase, we modified it to include fine-tuning on target tasks for a fair comparison. We utilize the open-source code of PEARL at https://github.com/katerakelly/oyster.
@@
PEARL+RND extends PEARL by incorporating RND to enhance exploration. Like SAC+RND, the ratio of extrinsic to intrinsic rewards is set to match SISL for fair comparison.
@@
SPiRL (Pertsch et al., 2021) is a skill-based RL algorithm that first learns a fixed low-level policy from an offline dataset and then trains a high-level policy for specific tasks. SPiRL’s loss function is detailed in Section 3, and for fair comparison, loss scaling is aligned with SISL. We utilize the open-source code of SPiRL at https://github.com/clvrai/spirl.
@@
SiMPL (Nam et al., 2022) is a skill-based meta-RL algorithm that uses both offline datasets and meta-train tasks. While it shares SISL’s approach of extracting reusable skills and performing meta-train and meta-test phases, SiMPL fixes the skill model without further updates during meta-training. SiMPL’s loss function is also detailed in Section 3, and SiMPL’s implementation uses the same hyperparameters as SISL to ensure a fair comparison. We utilize the open-source code of SiMPL at https://github.com/namsan96/SiMPL.
C.2 Environmental Details
@@
Each subtask involves manipulating one object to its target state, while a full task requires sequentially completing four subtasks. The agent earns a reward of 1 for each completed subtask, with a maximum score of 4 achievable within a 280-timestep horizon. For instance, an example task, illustrated in Fig. C.1, requires the agent to complete the sequence: microwave $\rightarrow$ kettle $\rightarrow$ bottom burner $\rightarrow$ slide cabinet. The observation space is a 60-dimensional continuous vector representing object positions and robot state information, while the action space is a 9-dimensional continuous vector. Based on the task setup from Nam et al. (2022), we expanded the meta-train task set by adding two additional tasks, resulting in a total of 25 meta-train tasks and 10 meta-test tasks. Detailed task configurations are provided in Table C.1.
Table C.1: List of meta-train tasks and meta-test tasks in Kitchen environment
@@
[…]
@@
An example task, depicted in Fig. C.2, requires the agent to sequentially complete: (shed2 $\rightarrow$ drawer), (eraser1 $\rightarrow$ container), and (pepsi2 $\rightarrow$ tray). The observation space is a 76-dimensional continuous vector, including object positions and robot state information, while the action space is an 8-dimensional continuous vector. The meta-train and meta-test sets include 25 and 10 tasks, respectively, similar to the configuration in the Kitchen environment. A detailed task list is provided in Table C.2.
Table C.2: List of meta-train tasks and meta-test tasks in Office environment
@@
Figure C.3: Maze2D: Visualization of simulation and meta-train/test tasks in Maze2D
@@
Fig. C.3 (a) provides a visualization of the Maze2D environment, while Fig. C.3 (b) illustrates the meta-train and meta-test tasks. In Fig. C.3 (b), green squares indicate the goal points for 40 meta-train tasks, and red squares represent the goal points for 10 meta-test tasks. All tasks share the same starting point at (10, 10), marked by a blue cross. The observation space is a 4-dimensional continuous vector containing the ball’s position and velocity, while the action space is a 2-dimensional continuous vector.
@@
Figure C.4: AntMaze: Visualization of simulation and meta-train/test tasks in AntMaze
@@
Fig. C.4 (a) shows a simulation image of the AntMaze environment, and Fig. C.4 (b) depicts the meta-train and meta-test tasks. In Fig. C.4 (b), green squares mark the goal points for 20 meta-train tasks, while red squares denote the goal points for 10 meta-test tasks. All tasks share a common starting point at (5, 5), indicated by a blue cross. The observation space is a 29-dimensional continuous vector that includes the Ant’s state and its $(x,y)$ coordinates, while the action space is an 8-dimensional continuous vector.
C.3 Construction of Offline Dataset
@@
C.4 Hyperparameter Setup
In this section, we outline the hyperparameter setup for the proposed SISL framework. For high-level policy training, we adopt the hyperparameters from SiMPL for the Kitchen and Maze2D environments. For the Office and AntMaze environments, we conduct hyperparameter sweeps using the Kitchen and Maze2D configurations as baselines.
To ensure a fair comparison, we inherit from SiMPL all hyperparameters that are shared with SISL, given its multiple loss functions, and we perform parameter sweeps only over SISL specific components, namely self-improving skill refinement and skill prioritization via maximum return relabeling. We explore prioritization temperature values $T\in[0.1,0.5,1.0,2.0]$ and KLD coefficients $\lambda^{\text{kld}}_{\text{imp}}\in[0,0.001,0.002,0.005]$ for skill exploration, selecting the best-performing configurations as defaults. Additionally, the ratio of intrinsic to extrinsic rewards is fixed at levels that show optimal performance in single-task SAC experiments.
For implementing the skill models ($\pi_{l},q,p$), we follow SPiRL by utilizing LSTM (Graves & Graves, 2012) for the skill encoder and MLP structures for the low-level skill policy and skill prior. For implementing the high-level models ($\pi_{h},Q_{h},q_{e}$), we follow SiMPL by utilizing Set Transformer (Lee et al., 2019) for the task encoder and MLP structures for the high-level policy and value function. Additionally, for implementing the SISL, we utilize MLP structures for $\pi_{\text{imp}}$, $Q_{\text{imp}}$, and $\hat{R}$. The detailed hidden network sizes are presented in Table C.3 and Table C.4. Table C.3 presents the network architectures (the number of nodes in fully connected layers) and the hyperparameters shared across all environments, while Table C.4 details the environment-specific hyperparameter setups.
Table C.3: Network Architecture and Shared Hyperparameters
Shared
Hyperparameters
Group Name Environments
@@ C.4 Hyperparameter Setup
High-level
Discount Factor $\gamma_{h}$ 0.99
@@
[…]
Low-level
Buffer size $\mathcal{B}^{i}_{\text{on}}$ 10K
@@
[…]
@@
Skill-Improvement
RND state dropout ratio 0.7
@@
Learning rate $\lambda^{\text{lr}}_{\text{imp}}$ 0.0003
[…]
Table C.4: Environmental Hyperparameters
Environmental
Hyperparameters
Group Name Environments
@@
High-level
Buffer size $\mathcal{B}^{i}_{h}$ 3000 3000 20000 20000
@@
Low-level
Skill refinement $K_{\text{iter}}$ 2000 2000 1000 2000
@@
Skill-Improvement
Buffer size $\mathcal{B}^{i}_{\text{imp}}$ 100K 200K 100K 300K
Discount factor $\gamma_{\text{imp}}$ 0.95 0.95 0.99 0.99
@@
Entropy coefficient $\lambda^{\text{ent}}_{\text{imp}}$ 0.2 0.2 0.1 0.1
KLD coefficient $\lambda^{\text{kld}}_{\text{imp}}$
0.005 (Expert)
0.001
0.001
0.001
@@
Appendix D Additional Comparison Results
In this section, we provide additional comparison results against baseline algorithms. Following Fig. 6, additional performance comparisons across all environments and noise levels are presented in Section D.1, limited offline dataset size in Section D.2, random noise injection in Section D.3, and diverse sub-optimal offline datasets in Section D.4.
D.1 Performance Comparison